find the area of the surface that is generated by revolving to portion of the curve y=x^3 between x=0 and x=1 about x axis?

1
Expert's answer
2022-04-11T16:04:29-0400

ANSWER "\frac{\pi}{27}\cdot (10\sqrt{10}-1)\cong3.5631"

EXPLANTION

If "f'(x)" is continuous on "[a,b]" and the curve "y=f(x)" is rotated about x-axis , the area of the surface of revolution so generated is

"S =2\pi \int_{a}^{b}\left | f(x) \right |\sqrt{1+\left ( f'(x) \right )^{2}}dx"

Let "f(x)=x^3" , "a=0, b=1" , then "f'(x)=3x^2" and

"S =2\pi \int_{0}^{1}x^3\sqrt{1+9x^4}dx"

Calculate the integral using the substitution "1+9x^4=t" . So, since "36x^3dx=dt" , we have

"S = \frac{2\pi}{36}\int_{0}^{1}36x^3\sqrt{1+9x^4}dx=\\= \frac{\pi}{18} \int_{1}^{10}\sqrt{t}dt=\frac{2\pi}{18\cdot 3}\cdot[t^{\frac{3}{2}}]_{1} ^{10} =\frac{2\pi}{18\cdot 3}\cdot10\sqrt{10} -\frac{2\pi}{18\cdot 3} = \frac{\pi}{27}\cdot (10\sqrt{10}-1)\cong3.5631"